Operation Summary for July 8, 2011 - July 11, 2011 |
The Integrated Luminosity for the period from 7/4/11 to 7/11/11 was 58.1 Pb-1 (inverse picobarns). NuMI reported receiving 3.55E18 protons on target during this same period. Friday July 8 The day shift began with Operations monitoring the Tevatron (TEV) store 8868 (88.48E30) and the Recycler stash (154.9E10), with the Antiproton Source (Pbar) stacking (18.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-1015 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, operators prepared MTA for access. NuMI personnel conducted a 30-minute target scan and then resumed normal running. Operations terminated store 8868 at 12:05 PM. Operators prepared the TEV, CDF, and D0 for access. TEV sector B4 vacuum work began. From 12:30 PM to 3:07 PM, the following jobs were conducted: Replaced the 400 MeV area Debuncher RF LCW pump, MI-60 South Anode power supply work, and TEV sector E48 kicker high voltage power supply work. CDF and D0 were out of access by 2:38 PM. At 3:58 PM, the P1 transfer line beam permit tripped off due to the voltage output of a power supply (HT702) going low; operators investigated. During the evening shift, Vacuum personnel reported at 5:27 PM that they had completed their sector B4 work and came out of the TEV. At 5:45 PM, the Linac RF quadrupole power supplies and other devices tripped off. Operators investigated and discovered a Linac line voltage drop; it recovered and everything reset. Cryo system personnel reported at 5:53 PM that the TEV B4 cryo system was cooling down. Operations disabled beam to T-1015 at 6 PM. Also at 6 PM, Operations turned off beam to NuMI; their target had developed a LCW leak and must be replaced. They estimated it would take two and a half weeks to replace the target. At 6:54 PM, a Booster RF station (BRF3) tripped off and continued to trip off for about an hour; operators bypassed the station, brought BRF19 on in its place, and called in a RF person to make repairs. Operations put the TEV through a dry squeeze, orbit check, and into shot setup at 11:19 PM. Saturday July 9 The midnight shift began with TEV in shot setup, with Operations monitoring the stash (7.3E10), with Pbar stacking (72.7E10), with NuMI off for target replacement, and with MiniBooNE taking beam. During the midnight shift, a Booster RF person stopped working on BRF3 at 12:14 AM; the station was left in bypass. Operations established store 8871 at 12:23 AM with an average initial luminosity (AIL) of 401.6E30. Operations established beam to T-1015 at 7:05 AM. During the day shift, at 2:01 PM, operators reported problems with MI beam extraction; operators found and corrected some ramp problems. Operations terminated store 8871 at 3:31 PM and put the TEV into shot setup. At 3:50 PM, a TEV BA cold compressor tripped off; operators notified the Cryo Coordinator. During the evening shift, Operations established store 8872 at 4:10 PM with and AIL of 361.4E30. Operations disabled beam to T-1015 at 6 PM. Operators reported that the MI South Anode power supply tripped off; it reset. Operators prepared NuMI for access at 8:13 PM. Sunday July 10 The midnight shift began with Operations monitoring store 8872 and the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8872 at 7:27 AM and put the TEV into shot setup. Operations established beam to T-1015 at 6:11 AM. Operators reported short downtimes due to a 750 KeV area Buncher trip, a Linac Klystron RF station (KRF3) spark, and to LRF1 & 3 and BRF14 trips. During the day shift, at 8:06 AM, operators reported MI coalescing problems; tuning didn’t help. Operators called in MI personnel to help. Operations put the TEV through a dry squeeze at 8:23 AM. Operators put Pbar in standby during the MI coalescing investigations. MI personnel reported that there was a problem with the MI damper system. Reboots didn’t work. Operators contacted a Controls person who also investigated. A Damper person came in and managed to get the damper system to restart at 11:23 AM. At 11:55 AM, the MiniBooNE horn tripped off due to high temperatures at the MI-12 service building. A MiniBooNE person said to leave the beam off until the temperatures came down. At 2:34 PM, the MI damper problems returned. MI personnel believed the damper was mis-phased. Attempts to restart and reboot didn’t work. While talking with Control personnel, everything started working again. During the evening shift, Operators put the TEV through a wet squeeze at 4:01 PM. Operations put the TEV into shot setup at 4:24 PM. The MI-12 service building temperatures came down and MiniBooNE resumed taking beam at 5:12 PM. Operations established store 8875 at 5:26 PM with an AIL of 395.6E30. D0 reported that their solenoid power supply tripped off at 5:47 PM due to their LCW pump tripping off; they brought their backup pump on line. Operations disabled beam to T-1015 at 6:10 PM. At 8:18 PM, the D0 solenoid tripped off again due to the backup LCW pump tripping off; they called in mechanical technicians to repair the pumps. Monday July 11 The midnight shift began with Operations monitoring store 8878, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, D0 reported that their LCW pump work had finished and they began taking data at 12:04 AM. At 12:12 AM, the I- Source vacuum pumps tripped off; operators reset the pumps. Operators reported at 12:53 AM, that the TEV sector A0 kicker room temperature alarmed high. Operators investigated and found that two of the three air conditioner (AC) units weren’t working; they contacted FESS personnel to investigate. At 1:05 AM, the MiniBooNE horn and beam permit tripped off. Operators contacted a MiniBooNE person who told them to wait a while and then try to reset it. The horn reset and beam returned at 1:30 AM. FESS personnel reported at 6:05 AM that they had gotten both AC units on and running for the A0 kicker room. Operations established beam to T-1015 at 4:23 AM. Operations terminated store 8875 at 6:19 AM and put the TEV into shot setup. At 6:40 AM, the MI interlock power supply loop dropped on a ground fault; it reset. Operators reported MI coalescing problems at 7:20 AM; they contacted MI personnel, but continued with shot setup. At 7:58 AM, operators reported that a lightning strike caused the loss of the stash and many other systems to trip off, including the Pelletron. The Plans for Monday The plans for today are to allow heat exchanger work around the TEV while beam is off. The Run Coordinator will allow the following jobs: Booster studies, a Linac Debuncher pump replacement, CDF and D0 access, and cryo system flywheel work. |
|
More Information Comments and Suggestions |
| last modified 07/11/2011 email Fermilab |
| Security, Privacy, Legal | |